OpenStudy (anonymous):

4^2(x + 5) - 11 = 245

OpenStudy (anonymous):

\[4^{2x + 5} - 11 = 245\]

OpenStudy (anonymous):

4^2x+5 = 256 256 = 2^8 = 4^4

OpenStudy (anonymous):

then all you have to do is to compare powers

OpenStudy (anonymous):

2(x + 5) = 4 x+5 = 2 x = -3